Find the following limits:



lim x approches 2 for x^2 + 4x -1/ x^2 -2x

1
Expert's answer
2021-11-03T17:13:58-0400

"\lim\limits_{x\rarr 2}\frac{x^2+4x-1}{x^2-2x}=?"

"\lim\limits_{x\rarr 2}\frac{x^2+4x-1}{x^2-2x}=\frac{(2)^2+4(2)-1}{2^2-2(2)}"

"=\frac{4+8-1}{4-4}"

"=\frac{11}{0}"

"=\infty"

"\therefore \lim\limits_{x\rarr 2}\frac{x^2+4x-1}{x^2-2x}=\infty"